# Graph-Enhanced EEG Foundation Model

### v2

@@ Abstract
Electroencephalography (EEG) signals provide critical insights for applications in disease diagnosis and healthcare. However, the scarcity of labeled EEG data poses a significant challenge. Foundation models offer a promising solution by leveraging large-scale unlabeled data through pre-training, enabling strong performance across diverse tasks. While both temporal dynamics and inter-channel relationships are vital for understanding EEG signals, existing EEG foundation models primarily focus on the former, overlooking the latter. To address this limitation, we propose a novel foundation model for EEG that integrates both temporal and inter-channel information. Our architecture combines Graph Neural Networks (GNNs), which effectively capture relational structures, with a masked autoencoder to enable efficient pre-training. We evaluated our approach using three downstream tasks and experimented with various GNN architectures. The results demonstrate that our proposed model, particularly when employing the GCN architecture with optimized configurations, consistently outperformed baseline methods across all tasks. These findings suggest that our model serves as a robust foundation model for EEG analysis.
@@ I Introduction
In this study, we propose a novel approach that integrates GNNs into a Transformer-based EEG foundation model to effectively capture inter-channel relationships. GNNs are particularly well-suited for modeling and representing complex relationships between entities, making them an ideal choice for capturing the intricate interactions among EEG channels. While the integration of GNNs with sequence modeling techniques, such as combining knowledge graphs (KGs) with LLMs, has been extensively explored in other fields, to the best of our knowledge, this study is the first to apply this combination to EEG analysis. This approach aims to establish a versatile and task-agnostic foundation model, offering significant potential for advancing EEG research and its applications.
@@ III Methods
Figure 1: Comparison between two model architectures during pre-training and fine tuning. Figures (a), (c) and (e) represent BENDR [5], while (b), (d) and (f) represent our proposed model.
@@ III-A BENDR
III-B Our Proposal
We extend the BENDR architecture described above by incorporating the learning of inter-channel relationships. Inter-channel relationships in EEG signals can be naturally represented using a graph structure, where each channel corresponds to a node and the edges represent the connectivity between channels. Graph Neural Networks (GNNs) are particularly well-suited for modeling such relationships in graph-structured data. Therefore, we propose integrating GNNs with BENDR to enhance its capability of learning inter-channel relationships.
@@ IV Experiments
We conducted experiments, including both pre-training and downstream tasks, under various conditions to evaluate the performance of our proposed model and gain insights into the research questions.
@@ IV-A2 Downstream Datasets
We evaluated our proposed model using the datasets from the following tasks.
@@ IV-B1 GNN architectures
We evaluated the performance of our proposed model by individually incorporating three standard GNN architectures: Graph Convolutional Networks (GCN) [7], Graph Attention Netowrks (GAT) [8] and GraphSAGE [9]. Each architecture was used consistently throughout a single set of pre-training and downstream tasks. All GNN implementations were based on PyTorch 2.3.1 and PyTorch Geometric 2.5.3.
@@ IV-C2 Differences of the Effect of GNNs Arising from Downstream Tasks (RQ2)
This section examines how the effect of introducing GNNs varies across downstream tasks by comparing the performance differences between the baseline and our model. Specifically, we evaluate our model configured as “GCN with edge weights”, identified as the most suitable architecture in the previous section, with the base model configuration set to “BENDR” and the sequence length adjuster implemented as a linear layer. For a fair comparison across tasks, we use the same configurations for the baseline. Hereafter, we refer to ‘our model configured as “GCN with edge weights”’ simply as ‘our model’.
As shown in Table II, on MMI, the baseline achieved a score of $0.646$, while our model achieved $0.849$, representing a $31.4\%$ improvement. On P300, the baseline achieved $0.568$, with our model improving performance by $8.53\%$. On ERN, the baseline achieved $0.522$, and our model showed a $3.11\%$ improvement. These results indicate that the higher the baseline performance on a task, the greater the relative improvement achieved by our model. Further analysis could investigate the relationship between model performance and task-specific characteristics, particularly the physiological features associated with each task, to gain deeper insights.
@@ IV-C4 Comparison of Base Model Configurations (RQ4)
This section examines whether incorporating GNNs is more effective when the base model configuration is “BENDR” or “Linear.” As shown in Tables II and III, when comparing the performance of models employing GNNs across all task-model combinations, the “BENDR” configuration outperformed the “Linear” configuration in 21 out of 24 cases. In contrast, among the baseline models, “Linear” exceeded “BENDR” in 2 out of 3 cases. Notably, all GNN-based models that outperformed the baseline were configured with “BENDR.” These results indicate that incorporating GNNs is more effective when the base model configuration is “BENDR.”
@@ V Conclusion
Foundation models for EEG analysis are particularly valuable due to the difficulty of collecting large amounts of labeled data and their ability to be applied to a wide range of EEG tasks with minimal computational time and cost. In this study, we propose an improved foundation model for EEG that leverages both inter-channel relationships and the temporal dynamics of EEG signals. The proposed architecture integrates GNNs, which are effective at learning relationships between entities, with a masked autoencoder-based framework. We evaluated the model using several GNN architectures across three downstream tasks. The results indicate that our proposed model, when employing the GCN architecture [7] with specific configurations, consistently outperformed the baseline across all tasks. These findings demonstrate that incorporating inter-channel relationships learning through GNNs enhances the model’s performance, establishing it as a more effective foundation model for EEG analysis.
As future work, we plan to evaluate the proposed model on a broader range of tasks to further demonstrate its versatility. Additionally, we aim to investigate the mechanisms underlying the observed improvements achieved through integrating graph structures, using techniques such as GNNExplainer [26]. And another potential direction is to expand our approach, which integrates inter-channel relationships and time-series information, to other base models or various self-supervised learning approaches Alternatively, we may design a novel architecture specifically optimized for EEG foundation models.
@@ VI Acknowledgments
This work is partially supported by JSPS KAKENHI Grant JP21K17749 and 23K28098.

### v3

# GEFM: Graph-Enhanced EEG Foundation Model

@@ Abstract
Electroencephalography (EEG) signals provide critical insights for applications in disease diagnosis and healthcare. However, the scarcity of labeled EEG data poses a significant challenge. Foundation models offer a promising solution by leveraging large-scale unlabeled data through pre-training, enabling strong performance across diverse tasks. While both temporal dynamics and inter-channel relationships are vital for understanding EEG signals, existing EEG foundation models primarily focus on the former, overlooking the latter. To address this limitation, we propose Graph-Enhanced EEG Foundation Model (GEFM), a novel foundation model for EEG that integrates both temporal and inter-channel information. Our architecture combines Graph Neural Networks (GNNs), which effectively capture relational structures, with a masked autoencoder to enable efficient pre-training. We evaluated our approach using three downstream tasks and experimented with various GNN architectures. The results demonstrate that our proposed model, particularly when employing the GCN architecture with optimized configurations, consistently outperformed baseline methods across all tasks. These findings suggest that our model serves as a robust foundation model for EEG analysis.
@@ I Introduction
In this study, we propose Graph-Enhanced EEG Foundation Model (GEFM), a novel approach that integrates GNNs into a Transformer-based EEG foundation model to effectively capture inter-channel relationships. GNNs are particularly well-suited for modeling and representing complex relationships between entities, making them an ideal choice for capturing the intricate interactions among EEG channels. While the integration of GNNs with sequence modeling techniques, such as combining knowledge graphs (KGs) with LLMs, has been extensively explored in other fields, to the best of our knowledge, this study is the first to apply this combination to EEG analysis. This approach aims to establish a versatile and task-agnostic foundation model, offering significant potential for advancing EEG research and its applications.
@@ III Methods
Figure 1: Comparison between two model architectures during pre-training and fine tuning. Figures (a), (c) and (e) represent BENDR [5], while (b), (d) and (f) represent our proposed model, GEFM.
@@ III-A BENDR
III-B Our Proposal: GEFM
We propose Graph-Enhanced EEG Foundation Model (GEFM), extending the BENDR architecture described above by incorporating the learning of inter-channel relationships. Inter-channel relationships in EEG signals can be naturally represented using a graph structure, where each channel corresponds to a node and the edges represent the connectivity between channels. Graph Neural Networks (GNNs) are particularly well-suited for modeling such relationships in graph-structured data. Therefore, we propose integrating GNNs with BENDR to enhance its capability of learning inter-channel relationships.
@@ IV Experiments
We conducted experiments, including both pre-training and downstream tasks, under various conditions to evaluate the performance of GEFM and gain insights into the research questions.
@@ IV-A2 Downstream Datasets
We evaluated GEFM using the datasets from the following tasks.
@@ IV-B1 GNN architectures
We evaluated the performance of GEFM by individually incorporating three standard GNN architectures: Graph Convolutional Networks (GCN) [7], Graph Attention Netowrks (GAT) [8] and GraphSAGE [9]. Each architecture was used consistently throughout a single set of pre-training and downstream tasks. All GNN implementations were based on PyTorch 2.3.1 and PyTorch Geometric 2.5.3.
@@ IV-C2 Differences of the Effect of GNNs Arising from Downstream Tasks (RQ2)
This section examines how the effect of introducing GNNs varies across downstream tasks by comparing the performance differences between the baseline and GEFM. Specifically, we evaluate GEFM configured as “GCN with edge weights”, identified as the most suitable architecture in the previous section, with the base model configuration set to “BENDR” and the sequence length adjuster implemented as a linear layer. For a fair comparison across tasks, we use the same configurations for the baseline. Hereafter, we refer to ‘GEFM configured as “GCN with edge weights”’ simply as ‘GEFM’.
As shown in Table II, on MMI, the baseline achieved a score of $0.646$, while GEFM achieved $0.849$, representing a $31.4\%$ improvement. On P300, the baseline achieved $0.568$, with GEFM improving performance by $8.53\%$. On ERN, the baseline achieved $0.522$, and GEFM showed a $3.11\%$ improvement. These results indicate that the higher the baseline performance on a task, the greater the relative improvement achieved by GEFM. Further analysis could investigate the relationship between model performance and task-specific characteristics, particularly the physiological features associated with each task, to gain deeper insights.
@@ IV-C4 Comparison of Base Model Configurations (RQ4)
This section discusses whether incorporating GNNs is more effective when the base model configuration is “BENDR” or “Linear.” As shown in Tables II and III, when comparing the performance of models employing GNNs across all task-model combinations, the “BENDR” configuration outperformed the “Linear” configuration in 21 out of 24 cases. In contrast, among the baseline models, “Linear” exceeded “BENDR” in 2 out of 3 cases. Notably, all GNN-based models that outperformed the baseline were configured with “BENDR.” These results indicate that incorporating GNNs is more effective when the base model configuration is “BENDR.”
@@ V Conclusion
Foundation models for EEG analysis are particularly valuable due to the difficulty of collecting large amounts of labeled data and their ability to be applied to a wide range of EEG tasks with minimal computational time and cost. In this study, we propose Graph-Enhanced EEG Foundation Model (GEFM), a novel foundation model for EEG that leverages both inter-channel relationships and the temporal dynamics of EEG signals. The proposed architecture integrates GNNs, which are effective at learning relationships between entities, with a masked autoencoder-based framework. We evaluated the model using several GNN architectures across three downstream tasks. The results indicate that GEFM, when employing the GCN architecture [7] with specific configurations, consistently outperformed the baseline across all tasks. These findings demonstrate that incorporating inter-channel relationships learning through GNNs enhances the model’s performance, establishing it as a more effective foundation model for EEG analysis.
As future work, we plan to evaluate GEFM on a broader range of tasks to further demonstrate its versatility. Additionally, we aim to investigate the mechanisms underlying the observed improvements achieved through integrating graph structures, using techniques such as GNNExplainer [26]. And another potential direction is to expand our approach, which integrates inter-channel relationships and time-series information, to other base models or various self-supervised learning approaches. Alternatively, we may design a novel architecture specifically optimized for EEG foundation models.
@@ VI Acknowledgments
This work is partially supported by JSPS KAKENHI Grant JP21K17749 and JP23K28098.